OpenStudy (anonymous):

Alex has 520 yards of fencing to encolose a returangular area. A retaungle that maxmises the enclosed area has a length of _ yards and a width of _ yards. The maxium area is _____ yards ?

OpenStudy (turingtest):

perimeter: p=2(w+L)=520 w=260-L A=wL=260L-L^2 at max dA/dL=260-2L=0 L=130 w=260-L=130 A(130)=130^2=16900m^2

OpenStudy (anonymous):

huh

OpenStudy (turingtest):

note that the maximum area is found when w=L i.e. a square maximizes the area of a rectangle.

OpenStudy (turingtest):

what math are you in kandice?

OpenStudy (anonymous):

college algebra

OpenStudy (anonymous):

I don't get how to get the maxium... thats where im stuck

OpenStudy (turingtest):

well in that case ignore the dA/dL part because that's calculus so look at it this way: We are told the perimeter is 520. The formula for perimeter is p=2w+2L=520. The only way for you to solve this is by recognizing that the area is maximized by a square, so L=w. Now the above becomes p=2w+2w=520=4w w=130 now just find the area A=wL=130(130)=16900

OpenStudy (anonymous):

so the max would be 16900

OpenStudy (turingtest):

There is no way I know of to find the maximum of a given function without calculus, except by recognizing a special case like this from prior knowledge.

OpenStudy (turingtest):

yes that is the maximum area possible

OpenStudy (turingtest):

when w=L i.e. a square

OpenStudy (turingtest):

by the way the maximum area is given in yards-squared, not yards as it says above.